OpenStudy (gamer56):

What are the x- and y-intercepts of the line? 2x – 3y = –12 A. x-intercept 6, y-intercept –4 B. x-intercept –4, y-intercept 6 C. x-intercept 4, y-intercept –6 D. x-intercept –6, y-intercept 4

Mehek (mehek14):

do you know how to solve for them?

OpenStudy (gamer56):

I think its D. Am i right ? @thomas5267

Mehek (mehek14):

to find x-int, substitute 0 for y to find y-int, substitute 0 for x

Mehek (mehek14):

yea it's D you're correct

OpenStudy (gamer56):

Is D. Right or no ?

Mehek (mehek14):

right

OpenStudy (gamer56):

YES ! :D

Mehek (mehek14):

\(\ddot\smile\)

OpenStudy (gamer56):

@Mehek14 Can you answer this please ? Select the equation of the line that passes through the point (2, –1) and has slope 3 in point-slope form. A. (x – 2) = 3(y + 1) B. (y – 1) = 3(x + 2) C. (y + 1) = 3(x – 2) D. (x + 2) = 3(y – 1)

Mehek (mehek14):

open it in a separate post to avoid lag

OpenStudy (er.mohd.amir):

y-y1=m(x-x1) where m is slope and (x1,y1) is point.

OpenStudy (gamer56):

what is the answer

OpenStudy (gamer56):

@Er.Mohd.AMIR

OpenStudy (er.mohd.amir):

option C